OpenStudy (anonymous):

the temperature of a metal increase from 4 degree C to T degree C which is related by the equation T=4(1X2)^x when it was heated for x second. find: A_ the temperature of the metal when it is heated for 9 seconds. B_ the time taken, in second, need to raise the temperature from 5 dgree C to 20 degree C

OpenStudy (anonymous):

Emm, ehat do you mean by (1X2) ?

OpenStudy (anonymous):

It state 1.2 the point is at the center

OpenStudy (anonymous):



OpenStudy (shane_b):

If it's:\[T=4(1*2)^x\]The basic formula would be:\[T_f=T_i*2^x\]Where Tf and Ti are the final and initial temperatures respectively and x is the time in seconds. Assuming that's correct... Part A:\[T_f=4*2^9=?\] Part B:\[20=5*2^x=?\]Just solve for x in that one.